# Variable coherence model for free-electron laser pulses

### v1

@@ III.2 Joint Probability Distributions
Figure 6 shows a series of joint probability distributions that describe the likelihood of generating a FLASH pulse with a particular number of sub-pulses in both the time and frequency domains. In the zero coherence width case of panel (a), we observe a weak, positive correlation between time and frequency, with the most likely number of sub-pulses being 4 in both domains–consistent with our findings in figure 5. Also, we observe that there is a slightly increased likelihood of generating more sub-pulses in the frequency domain than the time domain. As the coherence width is increased towards the bandwidth in panels (b-c), we see that the distribution begins to narrow along the main diagonal, indicating increasing correlation strength. Comparing the four panels, we also see that larger coherence widths result in fewer sub-pulses within both domains, which is again consistent with the results of figure 5. Put together, these results demonstrate that the number of sub-pulses in the time and frequency domains is mildly correlated for the FLASH parameters, particularly for small coherence widths. We observe qualitatively similar trends for the LCLS-I and LCLS-II parameters, with analogous joint probability distributions to figure 5 shown in appendix A. Notably, for the LCLS-I parameters, we find a much more considerable bias towards the frequency domain, indicating a strong likelihood that a pulse will have more sub-pulses in the frequency domain than in the time domain. For LCLS-II, the joint probability distribution is nearly symmetric and already concentrated towards mostly one or two sub-pulses in both the time and frequency domains, reflecting the reduced duration of pulses with respect to the average bandwidth.
@@
Figure 6: Joint probably distributions for the sub-pulse numbers in the time and frequency domains for VCM pulses with increasing coherence widths – see the titles. All panels are obtained with the FLASH set of parameters in table 1.
@@
The correlation analysis for the intensities of sub-pulses in the time versus frequency domain is less straightforward than for the sub-pulse number, as there is no obvious one-to-one mapping between sub-pulses in either domains – recall that many VCM pulses have a different number of time vs frequency sub-pulses. Instead, we consider all possible pairings of time-frequency sub-pulse intensities within each VCM pulse and show the corresponding joint probability distribution in figure 7. At zero coherence width in panel (a), the bulk of intensity pairs lie in regions of low intensity for both domains without exhibiting substantial correlation with one another. As the coherence width is increased to 0.25 eV in panel (b), the distribution begins to take on a more defined structure, with the bulk of intensity pairings still concentrated towards lower intensities in time but higher intensities in frequency. This observation is consistent with our understanding of the coherence width’s effect on a pulse’s structure: the time domain requires larger coherence widths for the sub-pulse intensities to reach their asymptotic value due to the appearance of lower-intensity pedestals that weigh down the average – see the discussion around figure 3. When the coherence width equals the average bandwidth in panel (c), we see that most of the sub-pulse intensities in frequency are concentrated around the Fourier-limited limit while we still have a broad distribution of intensities in time. As the coherence width is further increased beyond the bandwidth in panel (d), the sub-pulse intensities in time begin to approach their asymptotic Fourier-limited value. Similar to the sub-pulse number of figure 5, we generally observe a weak correlation between sub-pulses in the time vs frequency domains. Instead, the joint probability distributions highlight the different rate at which sub-pulses in each domain approach their Fourier-limited value as the coherence width is increased.
@@
Appendix A presents the joint probability distributions of sub-pulse intensities in the time versus frequency domains for the LCLS-I and LCLS-II cases. For the LCLS-I parameter set, we observe qualitatively similar results to the FLASH case from figure 7. On the other hand, the LCLS-II parameter set yields a very distinct behavior: irrespective of the coherence with, we observe a strong (anti-)correlation between the time and frequency sub-pulse intensities, which tend to cluster around their Fourier-limited asymptotic values. We attribute these results to the limited number of sub-pulses – mostly one or two per pulse – again, due to the reduced pulse duration compared to the average bandwidth.
Figure 7: Joint probability distributions for the sub-pulse intensities in the time and frequency domains for VCM pulses with increasing coherence widths – see titles. All panels are obtained with the FLASH set of parameters in table 1.
IV Absorption Calculations
@@
$\sigma(\omega)$ $=$ $-\frac{1}{\omega}\Im\left\{\frac{\mathcal{F}[\ddot{d}](\omega)}{\mathcal{F}[E](\omega)}\right\},$ (4)
$S(\omega)$ $=$ $-\frac{2}{\omega}\Im\left\{\mathcal{F}[\ddot{d}](\omega)\mathcal{F}[E](\omega)^{*}\right\},$ (5)
@@
Figure 8: (a) Atomic potential used in our absorption simulations. (b) Absorption cross-sections obtained from coherent cos4 pulses across a range of intensities. (c) Spectral intensity for the $4.65\times 10^{14}$ W/cm2 peak intensity coherent cos4 pulse (dotted curve) and matching average VCM pulse with zero coherence width (solid).
The presence of $\mathcal{F}[E]$ in the denominator of equation (4) makes the absorption cross-section from individual VCM pulses highly sensitive to low-intensity dips that are characteristic of small coherence widths. Instead, we consider the response function of equation (5) and show the results in figure 9. Panel (a) shows the results for minimally coherent pulses with zero coherence width. The thin transparent curves, obtained by averaging the response function over 250 VCM pulses, yield significant fluctuation between batches, reflecting the large shot-to-shot fluctuations in the pulses – see the discussion in section III. After averaging over 4000 pulses, we obtain a smooth average response function that is notably different from the result at large coherence width – compare the thick opaque and thin black curves. In particular, we see that the maximum in the response function at 50.31 eV is about 10% smaller for the zero-coherence width curve compared to the 10 eV reference. Panel (b) shows a similar analysis for a coherence width of 1.5 eV, slightly below the pulses’ average bandwidth. Compared to the zero coherence width case, we see that the 250 VCM pulse batches yield average response functions that are closer to the grand average, and that the grand average, which is now converged over as few as 1000 pulses, is itself closer to the large-coherence limit – again, compare the thick colored and thin black curves.
@@
Figure 9: (a)-(b) Average response functions of equation (5) for VCM pulses with the coherence width specified in the titles. In each panel, the opaque thick curve corresponds to the grand averages while transparent curves are batch averages over 250 VCM pulses only. For reference, the thin black curve corresponds to the average response function calculated with a coherence width of 10 eV – very close to the fully coherent Fourier limit. (c) Difference between the grand average response functions and reference average at 10 eV as a function of the VCM coherence width. Horizontal sections at 0 and 1.5 eV correspond to the difference between the thick opaque and thin black curves in panels (a,b). In all panels, the vertical dotted line at 50.31 eV marks the frequency at which the reference average response function is maximized.
@@
Figure 9 (c) shows the evolution of the difference between the average response functions and reference at 10 eV with increasing coherence width. Unsurprisingly, we observe the largest difference at zero coherence width, followed by an overall smooth convergence as the coherence width is increased. Already at $\approx 5$ eV coherence width, the average response function is only marginally different from the reference at 10 eV. Above 10 eV, we have observed that the average response function becomes virtually independent of the coherence width (not shown), which motivates our choice of using the 10 eV results as our reference. As we observed in panels (a,b), for each coherence width, the difference is largest around 50.31 eV (dotted line) and is associated with the gradual increase of the response function maximum with increasing coherence width. We attribute this to our observation from section III, that an increased coherence width leads to increasing the average peak intensity of VCM pulses, therefore strengthening the magnitude of the response function.
V Conclusion
@@ VI Acknowledgments
The authors thank James Cryan for enlightening discussions. This material is based upon work supported by the National Science Foundation under Grant No. PHY-2207656.
Data Availability
@@
Data underlying the results presented in this paper are not publicly available at this time, but may be obtained from the authors upon reasonable request.
Disclosures
The authors declare no conflicts of interest.
Appendix A Additional sub-pulse statistics
@@
Figure 10 shows the sub-pulse intensity distributions as a function of the coherence width for the LCLS-I parameter set. As mentioned in III A., the LCLS-I case demonstrates similar behavior to the FLASH case within both domains. In the frequency domain, we see a large initial spread of possible sub-pulse intensities that begins to increase towards an asymptotic value as the coherence width is increased. In the time domain, the low-coherence width intensities concentrate towards lower values, but also increase towards an asymptotic value for large coherence widths.
@@
Figure 10: Same as figure 3 for the LCLS-I parameter set of table 1: Sub-pulse intensity distributions as a function of the coherence width in the (a) frequency and (b) time domains.
@@
Contrary to both the FLASH and LCLS-I cases, the sub-pulse intensity distributions for the LCLS-II case shown in figure 11 have distinct behavior. In the frequency domain, we find that the most likely sub-pulse intensity is maximized at zero coherence width, and decreases towards an asymptotic value for larger coherence widths. In the time domain, we see an initial concentration towards lower intensities, and a gradual rise to an asymptotic intensity. This behavior in the time domain is reminiscent of the FLASH and LCLS-I cases, although the growth towards this asymptotic value for the LCLS-II case spans a smaller range of intensities in comparison.
@@
Figure 11: Same as figure 3 for the LCLS-II parameter set of table 1: Sub-pulse intensity distributions as a function of the coherence width in the (a) frequency and (b) time domains.
@@
The joint probability distributions discussed in section III B. are shown in figures 12 and  13 for the LCLS-I case. Again we obtain similar results to the FLASH parameter set for both the sub-pulse number and intensities. We see in figure 12 that the joint probability distribution for the sub-pulse number demonstrates weak positive correlation for low coherence widths with a considerable bias towards the frequency domain. Then, as the coherence width increases, the strength of correlation increases along with it, while the overall number of sub-pulses decreases. In figure 13, we find that the sub-pulse intensities between domains exhibit correlation in the same way that they did in the FLASH case, where we found that smaller intensities in time implied a greater likelihood of higher intensities in frequency, particularly for intermediate coherence widths.
@@
Figure 12: Same as figure 6 for the LCLS-I parameter set of table 1: Joint probability distributions for the sub-pulse number in both the frequency and time domains.
Figure 13: Same as figure 7 for the LCLS-I parameter set of table 1: Joint probability distributions for the sub-pulse intensity in both the frequency and time domains.
@@
The joint probability distributions for the LCLS-II case are shown in figures 14 and 15. The sub-pulse number distributions in figures 14 exhibit considerable positive correlation for all coherence widths shown, and they demonstrate a very slight bias towards the frequency domain. For the sub-pulse intensity distributions shown in figures 15, we find that the coherence width has relatively little effect on the sub-pulse intensities between domains, as the sub-pulse intensities tend to cluster very close to their asymptotic values regardless of the coherence width, as demonstrated in figure 11.
@@
Figure 14: Same as figure 6 for the LCLS-II parameter set of table 1: Joint probability distributions for the sub-pulse number in both the frequency and time domains.
Figure 15: Same as figure 7 for the LCLS-II parameter set of table 1: Joint probability distributions for the sub-pulse intensity in both the frequency and time domains.

### v2

@@ 3.2 Joint Probability Distributions
Figure 4 (a-d) shows a series of joint probability distributions that describe the likelihood of simulating a FLASH pulse with a particular number of sub-pulses in both the time and frequency domains. In the zero coherence width case of panel (a), we observe a weak, positive correlation between time and frequency, with the most likely number of sub-pulses being 4 in both domains – consistent with our findings in figure 3 (g,h). Importantly, the large spread in the joint sub-pulse-number probability distribution suggest that there is little correlation, if any, between the sub-pulses in the two domains. We attribute the slightly increased likelihood of generating more sub-pulses in the frequency domain than the time domain to our choices of a slightly higher prominence criterion in time in our sub-pulse detection – see Sec. 2.2. As the coherence width is increased towards the bandwidth in panels (b-c), we see that the distribution begins to narrow along the main diagonal, indicating increasing correlation strength. Comparing the four panels, we also see that larger coherence widths result in fewer sub-pulses within both domains, which is again consistent with the results of figure 3. Put together, these results demonstrate that the number of sub-pulses in the time and frequency domains is mildly correlated for the FLASH parameters, particularly for small coherence widths. We observe qualitatively similar trends in our simulations with the LCLS-I and LCLS-II parameters, with analogous joint probability distributions to figure 4 shown in the supplemental document. For LCLS-II, the joint probability distribution is nearly symmetric and already concentrated towards mostly one or two sub-pulses in both the time and frequency domains, reflecting the reduced duration of pulses with respect to the average bandwidth.
@@
Figure 4: Simulated joint probably distributions for the sub-pulse (a-d) numbers and (e-h) intensities in the time and frequency domains for VCM pulses with increasing coherence widths – see the titles. All panels are obtained with the FLASH set of parameters in table 1.
@@
The correlation analysis for the intensities of sub-pulses in the time versus frequency domain is less straightforward than for the sub-pulse number, as there is no obvious one-to-one mapping between sub-pulses in either domains – recall that many VCM pulses have a different number of time vs frequency sub-pulses. Instead, we consider all possible pairings of time-frequency sub-pulse intensities within each VCM pulse and show the corresponding joint probability distribution in figure 4 (e-h). At zero coherence width in panel (e), the bulk of intensity pairs lie in regions of low intensity for both domains without exhibiting substantial correlation with one another. As the coherence width is increased to 0.25 eV in panel (f), the distribution begins to take on a more defined structure, with the bulk of intensity pairings still concentrated towards lower intensities in time but higher intensities in frequency. This observation is consistent with our understanding of the coherence width’s effect on a pulse’s structure: the time domain requires larger coherence widths for the sub-pulse intensities to reach their asymptotic value due to the appearance of lower-intensity pedestals that weigh down the average – see the discussion around figure 2. When the coherence width equals the average bandwidth in panel (g), we see that most of the sub-pulse intensities in frequency are concentrated around the Fourier-limited limit while we still have a broad distribution of intensities in time. As the coherence width is further increased beyond the bandwidth in panel (h), the sub-pulse intensities in time begin to approach their asymptotic Fourier-limited value. Similar to the sub-pulse number in panels (a-d), we generally observe a weak correlation between sub-pulses in the time vs frequency domains. Instead, the joint probability distributions highlight the different rate at which sub-pulses in each domain approach their Fourier-limited value as the coherence width is increased.
@@
The supplemental document presents the simulated joint probability distributions of sub-pulse intensities in the time versus frequency domains for the LCLS-I and LCLS-II parameters. For the LCLS-I parameter set, we observe qualitatively similar results to the FLASH case from figure 4. On the other hand, the LCLS-II parameter set yields a very distinct behavior: irrespective of the coherence width, we observe a strong (anti-)correlation between the time and frequency sub-pulse intensities, which tend to cluster around their Fourier-limited asymptotic values. We attribute these results to the limited number of sub-pulses – mostly one or two per pulse – again, due to the reduced pulse duration compared to the average bandwidth.
4 Absorption Calculations
@@
$\sigma(\omega)=-\frac{1}{\omega}\Im\left\{\frac{\mathcal{F}[\ddot{d}](\omega)}{\mathcal{F}[E](\omega)}\right\}\qquad\text{and}\qquad S(\omega)=-\frac{2}{\omega}\Im\left\{\mathcal{F}[\ddot{d}](\omega)\mathcal{F}[E](\omega)^{*}\right\},$ (5)
@@
Figure 5: (a) Atomic potential used in our absorption simulations. (b) Absorption cross-sections obtained from coherent cos4 pulses across a range of intensities. (c) Spectral intensity for the $4.65\times 10^{14}$ W/cm2 peak intensity coherent cos4 pulse (dotted curve) and matching average VCM pulse with zero coherence width (solid). (d,e) Average response functions of equation (5) for VCM pulses with the coherence width specified in the titles. In each panel, the opaque thick curve corresponds to the grand averages while transparent curves are batch averages over 250 VCM pulses only. For reference, the thin black curve corresponds to the average response function calculated with a coherence width of 10 eV – very close to the fully coherent Fourier limit. (f) Difference between the grand average response functions and reference average at 10 eV as a function of the VCM coherence width. Horizontal sections at 0 and 1.5 eV correspond to the difference between the thick opaque and thin black curves in panels (d,e). In all panels, the vertical dotted line at 50.31 eV marks the frequency at which the reference average response function is maximized.
The presence of $\mathcal{F}[E]$ in the denominator of $\sigma$ in equation (5) makes the absorption cross-section from individual VCM pulses highly sensitive to low-intensity dips that are characteristic of small coherence widths. Instead, we consider the response function $S$ and show the results in figure 5 (d-f). Panel (d) shows the results for minimally coherent pulses with zero coherence width. The thin transparent curves, obtained by averaging the response function over 250 VCM pulses, yield significant fluctuation between batches, reflecting the large shot-to-shot fluctuations in the pulses – see the discussion in section 3. After averaging over 4000 pulses, we obtain a smooth average response function that is notably different from the result at large coherence width – compare the thick opaque and thin black curves. In particular, we see that the maximum in the response function at 50.31 eV is about 10% smaller for the zero-coherence width curve compared to the 10 eV reference. Panel (e) shows a similar analysis for a coherence width of 1.5 eV, slightly below the pulses’ average bandwidth. Compared to the zero coherence width case, we see that the 250 VCM pulse batches yield average response functions that are closer to the grand average, and that the grand average, which is now converged over as few as 1000 pulses, is itself closer to the large-coherence limit – again, compare the thick colored and thin black curves.
@@
Figure 5 (f) shows the evolution of the difference between the average response functions and reference at 10 eV with increasing coherence width. Unsurprisingly, we observe the largest difference at zero coherence width, followed by an overall smooth convergence as the coherence width is increased. Already at $\approx 5$ eV coherence width, the average response function is only marginally different from the reference at 10 eV. Above 10 eV, we have observed that the average response function becomes virtually independent of the coherence width (not shown), which motivates our choice of using the 10 eV results as our reference. As we observed in panels (d,e), for each coherence width, the difference is largest around 50.31 eV (dotted line) and is associated with the gradual increase of the response function maximum with increasing coherence width. We attribute this to our observation from section 3, that an increased coherence width leads to increasing the average peak intensity of VCM pulses, therefore strengthening the magnitude of the response function.
5 Conclusion
@@
Acknowledgment The authors thank James Cryan for enlightening discussions. This material is based upon work supported by the National Science Foundation under Grant No. PHY-2207656.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
@@
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]
Supplemental document Additional sub-pulse statistics for the LCLS-I and LCLS-II sets of parameters of table 1.
Supplemental document
Additional sub-pulse statistics
@@
Figure S1 shows the simulated sub-pulse intensity distributions as a function of the coherence width for the LCLS-I parameter set. As mentioned in section 3.1 of the main document, the LCLS-I case demonstrates similar behavior to the FLASH case within both domains. In the frequency domain, we see a large initial spread of possible sub-pulse intensities that begins to increase towards an asymptotic value as the coherence width is increased. In the time domain, the low-coherence width intensities concentrate towards lower values, but also increase towards an asymptotic value for large coherence widths.
@@
Figure S1: Same as figure 2 (main document) for the LCLS-I parameter set of table 1 (main document): simulated sub-pulse intensity distributions as a function of the coherence width in the (a) frequency and (b) time domains.
@@
Contrary to both the FLASH and LCLS-I cases, the simulated sub-pulse intensity distributions for the LCLS-II case shown in figure S2 have distinct behavior. In the frequency domain, we find that the most likely sub-pulse intensity is maximized at zero coherence width, and decreases towards an asymptotic value for larger coherence widths. In the time domain, we see an initial concentration towards lower intensities, and a gradual rise to an asymptotic intensity. This behavior in the time domain is reminiscent of the FLASH and LCLS-I cases, although the growth towards this asymptotic value for the LCLS-II case spans a smaller range of intensities in comparison.
@@
Figure S2: Same as figure 2 (main document) for the LCLS-II parameter set of table 1 (main document): simulated sub-pulse intensity distributions as a function of the coherence width in the (a) frequency and (b) time domains.
@@
The simulated joint probability distributions discussed in section 3.2 of the main document are shown in figure S3 for the LCLS-I case. Again we obtain similar results to the FLASH parameter set for both the sub-pulse number and intensities. We see in figure S3 (a-d) that the joint probability distribution for the sub-pulse number demonstrates weak positive correlation for low coherence widths. Here as well, we attribute the bias of the distribution center towards the frequency domain to our choice of a higher prominence criterion in time in our sub-pulse detection – see section 2.2 of the main document. As the coherence width increases, the strength of correlation increases along with it, while the overall number of sub-pulses decreases. In figure S3 (e-h), we find that the sub-pulse intensities between domains exhibit correlation in the same way that they did in the FLASH case, where we found that smaller intensities in time implied a greater likelihood of higher intensities in frequency, particularly for intermediate coherence widths.
@@
Figure S3: Same as figure 4 (main document) for the LCLS-I parameter set of table 1 (main document): simulated joint probability distributions for the sub-pulse number (a-d) and intensity (e-h) in both the frequency and time domains.
@@
The simulated joint probability distributions for the LCLS-II case are shown in figure S4. The sub-pulse number distributions in figure S4 (a-d) exhibit considerable positive correlation for all coherence widths shown. For the sub-pulse intensity distributions shown in  (e-h), we find that the coherence width has relatively little effect on the sub-pulse intensities between domains, as the sub-pulse intensities tend to cluster very close to their asymptotic values regardless of the coherence width, as demonstrated in figure S2.
@@
Figure S4: Same as figure 4 (main document) for the LCLS-II parameter set of table 1 (main document): simulated joint probability distributions for the sub-pulse number (a-d) and intensity (e-h) in both the frequency and time domains.